OpenStudy (4meisu):

4meisu The first term of a geometry sequence is 200 and the sum of the first 4 terms is 3248. What is the common ratio, and find the 10th term.

OpenStudy (anonymous):

first term a1 = 200 sum of first four term = 3248 which means a1 * r^3 = 3248 200 * r^3 = 3248 r^3 = 3248/200 r^3 = 406/25 r = 2.532 hence the 10th term will be a10 = a1 * r^9 = 200 * (r^3)^3 = 200 * (406/25)^3 = 856619.7

OpenStudy (cwrw238):

200 * r^4 - 1 ------ = 3248 r - 1